# Lambda Expected Shortfall

### v1

@@ 7 Conclusion
This paper introduces the $\Lambda$-Expected Shortfall ($\Lambda$-ES), a novel and theoretically grounded generalization of Expected Shortfall (ES) that robustly extends the $\Lambda$-VaR framework. We obtain an explicit representation of $\Lambda$-ES and verify that it satisfies several crucial properties as a desired counterpart to $\Lambda$-VaR. In particular, we show that $\Lambda$-ES is the smallest quasi-convex and law-invariant risk measure that dominates $\Lambda$-VaR. The dual representation of $\Lambda$-ES further connects it to established results on quasi-convex cash-subadditive risk measures. Our RU representation of $\Lambda$-ES provides useful insights for its potential applications to optimization problems. Practically, $\Lambda$-ES shares the advantages of $\Lambda$-VaR in the sense of flexible risk preferences and confidence levels, but more importantly, provides a useful alternative to $\Lambda$-VaR with additional benefits in risk management such as quasi-convexity and robustness.
The introduction of $\Lambda$-ES opens several promising avenues for future research. On the theoretical side, further investigation into its elicitability and connections to other axiomatically defined risk measures would lead to a deeper understanding of this new class of risk measures. Empirical studies applying $\Lambda$-ES in portfolio optimization, capital allocation, and stress testing with real-world financial data can be conducted for assessing its practical impact. Moreover, the development of efficient numerical algorithms for computing $\Lambda$-ES and solving $\Lambda$-ES-constrained optimization problems presents an important challenge for computational finance. $\Lambda$-ES holds substantial potential to enhance risk quantification and management practices, particularly in regulatory contexts that demand flexible yet robust risk measures.
@@ A.1 An algebraic formulation
A.2 A formulation based on the Rockafellar–Uryasev relation
Another possible formulation of $\Lambda$-ES is based on the RU relation in (3). Namely, for a decreasing function $\Lambda:\mathbb{R}\to(0,1)$, we may define the following candidate risk measure
@@ A.2 A formulation based on the Rockafellar–Uryasev relation
Here, we use infimum because the minimum may not exist in general. Clearly, $\rho_{\Lambda}$ is monotone in both $\Lambda$ and $X$, is law-invariant, and specializes to $\mathrm{ES}_{\alpha}$ when $\Lambda\equiv\alpha$ for $\alpha\in(0,1)$. Comparing (A.3) to Theorem 4, we are optimizing (22) over the subset $\{(a,x)\in\mathbb{R}^{2}:a=x\}$ so that the optimum is larger. We have $\mathrm{ES}_{\Lambda}\leqslant\rho_{\Lambda}$ so that $\rho_{\Lambda}$ also dominates $\mathrm{VaR}_{\Lambda}$. However, the following counterexample shows that $\rho_{\Lambda}$ defined in (A.3) is not quasi-convex in general, and is thus not an ideal candidate for $\Lambda$-ES.

### v2

@@ 7 Conclusion
This paper introduces the $\Lambda$-ES, a new and theoretically grounded generalization of ES that robustly extends the $\Lambda$-VaR framework. We obtain an explicit representation of $\Lambda$-ES and verify that it satisfies several crucial properties as a desired counterpart to $\Lambda$-VaR. In particular, we show that $\Lambda$-ES is the smallest quasi-convex and law-invariant risk measure that dominates $\Lambda$-VaR. The dual representation of $\Lambda$-ES further connects it to established results on quasi-convex cash-subadditive risk measures. The RU formula of $\Lambda$-ES provides useful insights for its potential applications to optimization problems. Practically, $\Lambda$-ES shares the advantages of $\Lambda$-VaR as a flexible model for risk assessment, and it has additional benefits in risk management such as quasi-convexity, $L^{1}$-continuity, and dual representation, thus sharing the advantages of ES over VaR.
@@ A.1 An algebraic formulation
A.2 A formulation based on the Rockafellar–Uryasev formula
Another possible formulation of $\Lambda$-ES is based on the RU formula in (5). Namely, for a decreasing function $\Lambda:\mathbb{R}\to(0,1)$, we may define the following candidate risk measure
@@ A.2 A formulation based on the Rockafellar–Uryasev formula
Here, we use infimum because the minimum may not exist in general. Clearly, $\rho_{\Lambda}$ is monotone in both $\Lambda$ and $X$, is law invariant, and specializes to $\mathrm{ES}_{\alpha}$ when $\Lambda\equiv\alpha$ for $\alpha\in(0,1)$. Comparing (A.3) to Theorem 4, we are optimizing (20) over the subset $\{(a,x)\in\mathbb{R}^{2}:a=x\}$ so that the optimum is larger. We have $\mathrm{ES}_{\Lambda}\leqslant\rho_{\Lambda}$ so that $\rho_{\Lambda}$ also dominates $\mathrm{VaR}_{\Lambda}$. However, the following counterexample shows that $\rho_{\Lambda}$ defined in (A.3) is not quasi-convex in general, and is thus not an ideal candidate for $\Lambda$-ES.